Hello, I have a problem I can't figure out. A litre of cream has 9.2% butterfat. How much skim milk containing 2% butterfat should be added to the cream to obtain a mixture with 6.4% butterfat. I'm not sure what to use as the variable. Please help.
Answer by checkley71(8403)

6.4=9.2x + 2(1-x) or 6.4 = 9.2x + 2 - 2x or 4.4=7.2x or x=4.4/7.2 or x=.6111
1-.6111=.3889
proof: 9.2 x .6111 = 5.62 2 x .3889 = .78 or 5.62+.78=6.4.